OpenStudy (anonymous):

The 600-mile trip to the lake took twice as long as the 240-mile trip to the ocean. If the speed to the ocean was 6 mph less than the speed to the lake, find both speeds and both times.

OpenStudy (anonymous):

The trip to the ocean took H hours. The trip to the lake took 2H hours. The ocean was 240 miles away. The lake was 600 miles away. The velocity of the trip to the ocean was 240/H miles per hour. To the lake, 600/2H = 300/H miles per hour. So if we know that the trip to the ocean was 6mph slower than the other one, we get Velocity ocean = Velocity lake - 6mph. 240/H = 300/H - 6 240 = 300 - 6H 6H = 60 H = 10 hours. Velocity to ocean = 24 mph. Velocity to lake = 30 mph.

OpenStudy (anonymous):

Thank you <3